Список публикаций

Д.В. Хмелёв

9 октября 2001

СТАТЬИ В РЕЦЕНЗИРУЕМЫХ ЖУРНАЛАХ

1. Хмелёв Д.В. Восстановление линейных индексных выражений для сведения класса программ к линейному//Журнал вычислительной математики и математической физики РАН, 1998, т.38, вып.3 (март), с.532–544, переведено в “Computational Mathematics and Mathematical Physics”, 1998, vol. 38, No.3, pp.513–525.

2. Хмелёв Д.В. Предельные теоремы для несимметричных транспортных сетей//Фунд. и прикл. математика, 2001, том 7, вып.4, c.1259–1266.

3. Оселедец В.И., Хмелёв Д.В. Стохастические транспортные сети и устойчивость динамических систем//Теория вероятностей и её приложения, 2001, том 46, вып.1, c.147–154.

4. Оселедец В.И., Хмелёв Д.В. Глобальная устойчивость бесконечных систем нелинейных дифференциальных уравнений и неоднородные счётные цепи Маркова. Пробл. передачи информации, 2000, том 36, вып.1, с.60– 76, переведено в“Problems of Information Transmission”, 2000, vol.36, no.1, pp.54–70

5. Хмелёв Д.В. Распознавание автора текста с использованием цепей А.А. Маркова. Вестник МГУ, сер.9: филология, N2, 2000, с.115–126.

6. Khmelev D.V. Disputed Authorship Resolution through Using Relative Empirical Entropy for Markov Chains of Letters in Human Language Text//Journal of Quantitative Linguistics, 2000, Volume 7, No 3, pp.201–207.

7. Кукушкина О.В., Поликарпов А.А., Хмелёв Д.В. Определение авторства текста с использованием буквенной и грамматической информации//Проблемы передачи информации, 2001, т.37, вып.2, с.96–108. Translated in “Problems of Information Transmission”, pp. 172–184.

8. Khmelev D., Tweedie F. Using Markov Chains for Identification of Writers//Literary and Linguistic Computing, 2001, vol.16, no.4, pp.299–307.

9. Khanin K., Khmelev D., Rybko A., Vladimirov A. Steady solutions for FIFO networks//Moscow Mathematical Journal, vol.1, no.3, pp.407–419.

ТЕЗИСЫ КОНФЕРЕНЦИЙ И ПРЕПРИНТЫ

1. Afanassieva L.G., Khmelev D.V. Large transportation networks with finitedimensional state space. Asymptotic approach//Proc. of 7th Vilnus Conference on Math. Stat., 1998.

2. Афанасьева Л.Г., Хмелёв Д.В. Большие транспортные сети с конечномерным пространством состояний//Теория вероятностей и её приложения, 1999, том 44, вып. 1, с.157–158, переведено в “Theory of Probability and its Applications” 1999, vol.44, N1, p.142–143.

3. Khmelev D.V., Oseledets V.I. Mean-field approximation for stochastic transportation network and stability of dynamical system//Technical Report 445, University of Bremen, June 1999. Preprint.

4. Khmelev D., Spodarev E. Stability of infinite closed Jackson networks//XX International Seminar on Stability Problems for Stochastic Models, Lublin – Naleczow, 5-11 September, 1999, Abstracts, p.85.

5. Khmelev D., Polikarpov A.A. Basic Assumptions about a Sign’s Life Cycle for Mathematical Modelling of Language System Evolution//Proceedings of the fourth conference of the International Quantitative Linguistics Association edited by R.H. Baayen. Prague, August 24-26, 2000.

6. Khmelev D. Disputed Authorship Resolution Using Relative Empirical Entropy For Markov Chain of Letters in a Text//Proceedings of the fourth conference of the International Quantitative Linguistics Association edited by R.H. Baayen. Prague, August 24-26, 2000.

7. Khmelev D.V., Spodarev E. Infinite closed Jackson networks. Technical Report Math/Inf/00/07, Friedrich-Schiller-Universität Jena, April 2000.

8. Khmelev D., Spodarev E. “Non-homogeneous zero-range interaction processes”, resume of the contribution to the XX Seminar on Stability Problems for Stochastic Models, Lublin - Naleczow, september 5-11 1999.

9. Хмелёв Д.В. Сложностной подход к задаче определения авторства//Тезисы конгресса �Русский язык: исторические судьбы и современность�, филологический факультет МГУ, Москва, март 2001, с.426-427.

В ПОДГОТОВКЕ

1. Khanin K., Khmelev D. Homeomorphisms of the circle with the break-type singularity I. Periodic Rotation Numbers. (in preparation, 2001)

2. Khmelev D. Periodic trajectories for homeomorphisms of the circle with two break-type singularities. (in preparation, 2001)

НЕОПУБЛИКОВАННЫЕ МАНУСКРИПТЫ

1. Хмелёв Д.В. О программе “Репетитор”“Самиздат”//http://zhurnal. lib.ru/h/hmelew_d_w/